OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2

OpenStudy (diyadiya):

Ok lets re-write the first equation 2x+y=1 subtract 2x from both sides y=1-2x Substitute this value of y in the second equation 4x+2y=-2 4x+ 2(1-2x)=-2 Now find x

OpenStudy (anonymous):

Pick your favorite of the two equations and your favorite of the two variables. Solve that equation for that variable. Going now to the OTHER equation, substitute in for that variable and solve.

OpenStudy (anonymous):

As Diddy has so kindly demonstrated.

OpenStudy (anonymous):

Thank you guys .

OpenStudy (diyadiya):

@Kaylaaa understood ?

OpenStudy (anonymous):

Yes , thankk youuu.

OpenStudy (diyadiya):

You're Welcome :)

OpenStudy (anonymous):

My pleasure =) Once you have the one variable solved, it's easy to substitute into either of the first equations and find the other variable.